• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 23:05:14 on May20,2008

    Entry number 239000

    This entry is a followup to: 238993

    keyword=TS11 fixed finally

    Bill Gunning swapped the power supply, Bob Michaels checked DAQ from the hall and everything is working fine. While they're closing the hall to go to the beam permit, cosmic has been started from the counting house to check the DAQ finally.